[Ncep.nhc.nco_contacts] TSB Morning Rounds - Mon Oct 24, 2016

TSB Admin nhc.tsbadmin at noaa.gov
Mon Oct 24 12:46:36 UTC 2016

Here is today's summary of NHC computer operations:
--- Mark


There was a problem with the 10/23/0000 surface analysis that was most  
likely due to a problem with one of NHC's N-AWIPS system. Mark will follow  
up with Andrew and Mike T. to get more details.
The web browser on the Pacific ATCF was not connecting to the internet or  
logging on to the HSU gmail account. Dave did some cleanup this morning and  
it appear to be working properly now.

The HWRF tracker for 91E did not work properly even though the fields  
showed a well-defined TC. The tracker did try to run since there was a line  
in the A-deck, but it had all zeros. This was most likely due to limitation  
of the tracker software itself. Mark let EMC know of the problem.

EP genesis area 37 wasn't available in the GUI when HSU entered  
probabilities at 00Z on Oct 22nd. They manually entered them into the  
e-deck and everything looks fine for 06Z. TSB will investigate.
Other items:
The problems we had last Friday reaching some external websites such as  
twitter were due to a nationwide denial of service attack.

Work on the media center continues this week.
RFC 192 to be implemented Tue 10/25 to turn off TAFB experimental rainfall  

RFC 205 to be implemented Tue 10/25 to send GOES WV mosaic to WCOSS so the  
TCGI program will work for potential TCs near the coast of Africa.

-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20161024/8e5ff187/attachment.html 

More information about the Ncep.nhc.nco_contacts mailing list